OpenStudy (anonymous):

How to take the derivative of f(x)= square root x+2, x=7

OpenStudy (anonymous):

\[\sqrt{x+2}\]?

OpenStudy (anonymous):

\[\sqrt{x+2}\] x= 7

OpenStudy (anonymous):

\[\sqrt{x+2}=(x+2)^{1/2}\] power rule / chain rule gets \[dy/dx=(1/2)(x+2)^{-1/2}=1/(2\sqrt{x+2})\]

OpenStudy (anonymous):

sub in 7 to get \[1/2\sqrt{9}\]= 1/(2x3) = 1/6

OpenStudy (anonymous):

Thank you it has help me alot